OpenStudy (anonymous):

Simpify: (tg^2x-sin^2x)/(ctg^2x-cos^2x)

OpenStudy (anonymous):

tan^6(x)

OpenStudy (anonymous):

i need steps

OpenStudy (anonymous):

I am not good with that, you might want to get someone else for that

OpenStudy (anonymous):

okey

OpenStudy (shayaan_mustafa):

Hi rolands111 :) Can you write your question as it is easy to parse?

OpenStudy (anonymous):

How to get tan^6(x) from this: (tg^2x-sin^2x)/(ctg^2x-cos^2x)

OpenStudy (shayaan_mustafa):

wait I am asking you to wirte in standard form. what are t and c in you quesiton?

OpenStudy (shayaan_mustafa):

\[\Large \frac{\tan^2x-\sin^2x}{\cot^2-\cos^2x}\]Is this your question look like?

OpenStudy (anonymous):

yeah

OpenStudy (shayaan_mustafa):

Are you there friend?

OpenStudy (shayaan_mustafa):

Write tanx and cotx in the form of sinx and cosxDid you get it?